The treasurer of a large corporation wants to invest $14 million in excess short-term cash in a particular money market investment. The prospectus quotes the instrument at a true yield of 6.46 percent; that is, the EAR for this investment is 6.46 percent. However, the treasurer wants to know the money market yield on this instrument to make it comparable to the T-bills and CDs she has already bought. If the term of the instrument is 92 days, what are the bond equivalent and discount yields on this investment?
Answer #1

HPY=(1+EAY)^(t/365)-1=(1+6.46%)^(92/365)-1=1.5904%

HPY=FV/P-1
=>P=FV/(1+HPY)=1000/(1+1.5904%)=984.3449775

Discount Yield=(1-P/F)*360/t=(1-984.3449775/1000)*360/92=6.1259%

Money Market yield=HPY*360/t=1.5904%*360/92=6.2233%